# Notions of enriched purity

### v2

@@ 1. Introduction
The concept of purity originated in the theory of abelian groups ([23]) and it is now well established both for modules ([22]) and in model theory in general ([29]). The monograph [2] made purity a key concept of the theory of accessible categories. For instance, given a locally $\lambda$-presentable category $\mathcal{K}$, it makes possible to characterize classes of objects, injective with respect to a family of morphisms whose domains and codomains are $\lambda$-presentable, as classes closed under products, $\lambda$-directed colimits and $\lambda$-pure subobjects (see [27]). Model-theoretically, this is a characterization of classes of models axiomatizable by regular theories in the logic $L_{\lambda\lambda}$ of $\lambda$-ary formulas.
@@
where $A$ and $B$ are $\lambda$-presentable, there is a morphism $t\colon B\to K$, such that $tg=u$ (see [2, 2.27]). Equivalently, given $u\colon A\to K$ then the existence of $v\colon B\to L$ such that $vg=fu$, implies the existence of $t\colon B\to K$ with $tg=u$. Since the existence of a morphism can be reformulated using surjectivity in the category $\operatorname{\bf Set}$ of sets, ordinary purity stems from the class $\mathcal{E}$ of surjections in $\operatorname{\bf Set}$. Building on [15], our enriched purity will depend on a chosen class $\mathcal{E}$ of morphisms in $\mathcal{V}$ which, in addition, arises as the left class of a factorization system. In metric-enriched categories these are the dense morphisms.
@@
For $\mathcal{V}=\operatorname{\bf Met},\operatorname{\bf CMet},\operatorname{\bf Ban}$ or $\omega$-$\operatorname{\bf CPO}$ with $\mathcal{E}$ consisting of the dense maps, then $\mathcal{G}=\{I\}$ is the singleton determined by the monoidal unit (so that powers by $\mathcal{G}$ are trivial); instead, when $\mathcal{V}=\mathbf{DGAb}$ and $\mathcal{E}$ is formed by the regular epimorphisms, then $\mathcal{G}=\{P_{n}\}_{n\in\mathbb{Z}}$ is given by the suspensions and desuspensions of the mapping cone of the unit (Example 8.4).
@@
While in this paper we already answer important questions about the relevance of enriched purity, there are still many open problems that wait to be addressed. For instance, when working with categories of structures of some relational language, ordinary pure morphisms correspond to the model theoretical notion of morphisms elementary with respect to positive-primitive formulas [2, 5.15]. A first glance at an enriched counterpart is given in Section 9. However, a general answer cannot be provided at this stage for the simple reason that enriched counterparts of these logical concepts have not yet been defined. We believe that a thorough study of enriched purity will be essential for the development of enriched notions of languages and models, thus covering an outstanding gap in the literature.
@@ 2. Background
Given a $\mathcal{V}$-category $\mathcal{K}$ and an ordinary functor $H\colon\mathcal{C}\to\mathcal{K}_{0}$ into the underlying ordinary category of $\mathcal{K}$, which then corresponds to a $\mathcal{V}$-functor $\mathcal{C}_{\mathcal{V}}\to\mathcal{K}$ from the free $\mathcal{V}$-category on $\mathcal{C}$. When $\mathcal{C}$ is small, the (conical) limit of $H$ in $\mathcal{K}$ is the data of an object $\lim H\in\mathcal{K}$ together with isomorphisms in $\mathcal{V}$
@@
that are $\mathcal{V}$-natural in $K\in\mathcal{K}$, where on the right-hand-side we have the ordinary limit in $\mathcal{V}_{0}$. It follows that $\lim H$, when it exists, is also the limit of $H$ in $\mathcal{K}_{0}$; moreover $\lim H$ corresponds to the weighted limit of $H$ weighted by $\Delta I\colon\mathcal{C}_{\mathcal{V}}\to\mathcal{V}$ (see [12, 3.51]).
@@
Local presentability will also be a key concept in the following sections. For ordinary locally presentable categories our standard reference is [2]. In the enriched setting the concept was introduced by Kelly [13]; given a $\mathcal{V}$-category $\mathcal{K}$ with (conical) $\lambda$-filtered colimits, we say that $K\in\mathcal{K}$ is $\lambda$-presentable (in the enriched sense) if $\mathcal{K}(K,-)\colon\mathcal{K}\to\mathcal{V}$ preserves $\lambda$-filtered colimits. We denote by $\mathcal{K}_{\lambda}$ the full subcategory of $\mathcal{K}$ spanned by the $\lambda$-presentable objects; this is closed in $\mathcal{K}$ under $\lambda$-small conical colimits and copowers by objects in $\mathcal{V}_{\lambda}$.
@@
For the remainder of this paper we shall assume our base of enrichment $\mathcal{V}$ to be locally $\lambda$-presentable as a closed category [13], meaning that $\mathcal{V}_{0}$ is locally $\lambda$-presentable, $I$ is $\lambda$-presentable and the $\lambda$-presentables are closed under tensor product (in particular $\mathcal{V}$ is then locally $\lambda$-presentable as a $\mathcal{V}$-category). In that context, many results from the ordinary theory of local presentability extend to the enriched context; see [13] for details.
Finally, (orthogonal) factorization systems will be central in the definition of our enriched notions of purity. Following [9] we will say that a factorization system $(\mathcal{E},\mathcal{M})$ on a category $\mathcal{K}$ is proper if every element of $\mathcal{E}$ is an epimorphism and every element of $\mathcal{M}$ a monomorphism.
@@ 3. Enriched purity
Let us introduce some notation concerning the maps involved in our (first) enriched notion of purity:
@@
of the map $\mathcal{K}(g,L)$. Let then $\mathcal{P}(g,f)$ be the pullback of $\mathcal{P}(g,L)$ along $\mathcal{K}(A,f)$, together with the induced map $r\colon\mathcal{K}(B,K)\to\mathcal{P}(g,f)$ as depicted below.
@@
We will say that $f\colon K\to L$ is $\mathcal{E}$-pure with respect to $g$ if the map $r$ above is in $\mathcal{E}$. We say that $f$ is $(\lambda,\mathcal{E})$-pure if it is $\mathcal{E}$-pure with respect to every $g\colon A\to B$ with $A$ and $B$ $\lambda$-presentable.
Since the elements of $\mathcal{M}$ are stable under pullback, the map $r$ is $\mathcal{E}$ if and only if $\mathcal{P}(g,f)$ is the $(\mathcal{E},\mathcal{M})$-factorization of $\mathcal{K}(g,K)$. Thus $f\colon K\to L$ is $(\lambda,\mathcal{E})$-pure if and only if for every $g\colon A\to B$, with $A$ and $B$ $\lambda$-presentable, the $(\mathcal{E},\mathcal{M})$-factorization of $\mathcal{K}(g,K)$ is obtained by pulling back the $(\mathcal{E},\mathcal{M})$-factorization of $\mathcal{K}(g,L)$.
@@
Following [25], a morphism $f\colon V\to W$ in $\mathcal{V}$ is called surjective if $\mathcal{V}_{0}(I,f)$ is a surjection of sets; a map is called injective if it satisfies the unique right lifting property with respect to surjections. Quite often, (surjective, injective) forms a proper factorization system on $\mathcal{V}$ (see [25]).
@@
    Consider $\mathcal{V}$ such that (surjective, injective) is a factorization system, and note that by [25, 3.4], injective maps are monomorphisms. Following Notation 3.1, it follows that $\mathcal{V}_{0}(I,\mathcal{P}(g,L))$ is the (epi, mono) factorization of $\mathcal{K}_{0}(g,L)$, so that
@@
    Given $\mathcal{V}$, consider the factorization system $(\mathcal{V}^{\to},\mathcal{V}^{\cong})$, where $\mathcal{V}^{\cong}$ is the class of all isomorphisms and $\mathcal{V}^{\to}$ the class of all morphisms. Then, every morphism $f$ is $\mathcal{V}^{\to}$-pure with respect to any $g\colon A\to B$ since the map $r$ of Notation 3.1 always lies in $\mathcal{V}^{\to}$.
@@
    Conversely, consider the factorization system $(\mathcal{V}^{\cong},\mathcal{V}^{\to})$ on $\mathcal{V}$. Given $f\colon K\to L$ and $g\colon A\to B$, it follows that $\mathcal{P}(g,L)=\mathcal{K}(B,L)$ and hence $\mathcal{P}(g,f)=\mathcal{K}^{\to}(g,f)$ is the hom in the $\mathcal{V}$-category of arrows $\mathcal{K}^{\to}$. Then $f$ is $\mathcal{V}^{\cong}$-pure with respect to $g$ if and only if $\mathcal{K}(B,K)\cong\mathcal{K}^{\to}(g,f)$, so that in particular for every pair $u\colon A\to K$ and $v\colon B\to L$, such that $fu=vg$, there is a unique $t\colon B\to K$ such that $ft=v$ and $tg=u$. It is easy to see that, for $\mathcal{K}$ locally $\lambda$-presentable, $(\lambda,\mathcal{V}^{\cong})$-pure morphisms are just the isomorphisms. On one hand, isomorphisms are $\mathcal{V}^{\cong}$-pure with respect to any morphism. Conversely assume that $f$ is $(\lambda,\mathcal{V}^{\cong})$-pure, and write $f$ as a $\lambda$-directed colimit of morphisms $f_{i}\colon K_{i}\to L_{i}$ between $\lambda$-presentable objects (see [2, 1.55(1)]); then the $\mathcal{V}^{\cong}$-purity property of $f$ with respect to the $f_{i}$’s, implies that $f$ can also be written as the colimit of the maps $1\colon L_{i}\to L_{i}$, showing that it is an isomorphism.
@@
    Let $\operatorname{\bf CMet}$ be the category of complete metric spaces and nonexpanding maps, where we allow distances $\infty$. Here, we have a proper factorization system (dense, isometry) (see [3, 3.16(2)]). Under this factorization system, given $g$ and $f$ as above, $\mathcal{P}(f,g)$ consists of those $u\colon A\to K$ such that for every $\varepsilon>0$ there exists $v\colon B\to L$ for which $fu\sim_{\varepsilon}vg$ (meaning that $d(fu,vg)\leq\varepsilon$). Then $f$ is $\mathcal{E}$-pure with respect to $g$ if and only if for any $u\in\mathcal{P}(f,g)$ as above and for any $\varepsilon>0$, there exists $t\colon B\to K$ such that $tg\sim_{\varepsilon}u$. A priori this doesn’t correspond to a notion of purity studied for metric-enriched categories. However it is easy to see that every $\lambda$-ap-pure morphism (as defined in [28]) is $(\lambda,\mathcal{E})$-pure; the converse is also true, making the two notions equivalent, and will be proved in Section 6.
@@
    In the category $\operatorname{\bf Met}$ of metric spaces and nonexpanding maps, we have two factorization systems: (dense, closed isometry) and (surjective, isometry) (see [3, 3.16]). The first one yields the same concept of purity described in (4) above, while the second one gives the usual (unenriched) concept of purity, as in (1). Observe that surjections in $\operatorname{\bf CMet}$ are the usual surjective maps but (surjective, injective), as defined in Remark 3.3, does not form a factorization system. For instance the morphism $\mathbb{N}\to\{\frac{1}{n}|n\in\mathbb{N}\}\cup\{0\}$ (sending $n$ to $\frac{1}{n}$), where the domain is discrete and the codomain has the usual metric, does not factorize.
@@
    Let $\operatorname{\bf MGra}$ be the category of oriented multigraphs with loops, i.e., the presheaf category on the category of chains $1$ and $2$. We have the factorization systems (surjective on vertices, injective on vertices) and (surjective on vertices and edges, injective on vertices and edges). The first one coincides with that from 3.3 and gives the usual (unenriched) concept of purity while the second one yields a seemingly stronger concept. Vertices of $\operatorname{\bf MGra}(A,B)$ correspond to morphisms $V\to\operatorname{\bf MGra}(A,B)$ (where $V$ is the terminal graph with one vertex and one edge); that is, to morphisms $A\cong V\times A\to B$. Edges of $\operatorname{\bf MGra}(A,B)$ correspond to morphisms $E\to\operatorname{\bf MGra}(A,B)$ (where $E$ the multigraph with two vertices and one edge between each pair of vertices); that is, to morphisms $E\times A\to B$ which can be interpreted as homotopies between morphisms. Now, purity with respect to (surjective on vertices and edges, injective on vertices and edges) says that, in addition to the ordinary lifting property, we also have an homotopy lifting property: every homotopy $u\sim u^{\prime}$ which extends to a homotopy $(u,v)\sim(u^{\prime},v^{\prime})$ can be lifted to a homotopy $t\sim t^{\prime}$ for which $tg\sim t^{\prime}g$ is $u\sim u^{\prime}$. But this follows from the ordinary concept of purity applied to
    $E\times A$$E\times B$$K$$L$$E\times g$$\bar{v}$$f$$\bar{u}$
    where $\bar{u}$ corresponds to $u\sim u^{\prime}$ an $\bar{v}$ to $v\sim v^{\prime}$. Thus $(\lambda,\mathcal{E})$-pure morphisms with respect to this factorization system coincides with ordinary $\lambda$-pure morphisms.
  7. (7)
    In the category $\operatorname{\bf SSet}$ of simplicial sets, we have a sequence of factorization systems $(\mathcal{E}_{n},\mathcal{M}_{n})$ where $\mathcal{E}_{0}$ are mappings surjective on vertices, $\mathcal{E}_{1}$ surjective on edges, etc. The corresponding purities start with the usual purity and successively have $n$-homotopy lifting properties. But, again, all these purities coincide.
  8. (8)
    Let $\mathcal{V}=\mathbf{Gra}$ be the cartesian closed category of directed graphs (possibly) without loops. Consider the (regular epi, mono) factorization system. Given $f\colon K\to L$, this is pure in the enriched sense if for any morphism $g\colon A\to B$ between finitely presentable objects, and any vertices $u\in\mathcal{K}(A,K)$ and $v\in\mathcal{K}(B,L)$ for which $vg=fu$, there exists a vertex $t\in\mathcal{K}(B,K)$ such that $tg=u$. Since not every vertex is a morphism in the underlying category (only the vertices with a loop are such), the notion of enriched purity is not the ordinary one on the underlying category; this is essentially because the unit is not regular projective. Similarly, surjections from 3.3 do not yield a factorization system (see [25, 2.7(3)]).
@@
    $\mathcal{V}=\mathbf{Ban}$. We can consider the (dense, isometry) factorization system, which induces a notion of purity similar to that of $\mathbf{Met}$ — see in particular Examples 3.21 and 4.20. There is also the (strong epi, mono) factorization system; a strong epimorphism in $\mathbf{Ban}$ is a morphism which is surjective on the unit ball. Since $I=\mathbb{C}$, these are surjections those from 3.3 ($e\in\mathbf{Ban}$ is a strong epi if and only if $\mathbf{Ban}_{0}(I,e)$ is surjective). Thus the notion of purity corresponding to this coincides with the ordinary one on the underlying category.
  11. (11)
    $\mathcal{V}=\mathbf{wSet}$ the category of weighted sets [10]. An object is a set $X$ together with a function $|-|\colon X\to[0,\infty]$ called weight; a morphism $f\colon X\to Y$ is a function such that $|f(x)|\leq|x|$ for any $x\in X$. The (regular epi, mono) factorization system induces a notion of purity which asks the lifing property for morphisms $u\in\mathcal{K}(A,K)$ and $v\in\mathcal{K}(B,L)$ which might have non-zero weight, while in ordinary purity we always need to assume $|u|=|v|=0$. However the two notions coincide when $\mathcal{K}$ has copowers (same argument used for graphs and simplicial sets).
@@
the squares $(a)$, $(b)$, and $(a+b)$ are pullbacks. By $\mathcal{E}$-purity of $f_{2}$, the map $r_{2}$ lies in $\mathcal{E}$ and thus $\mathcal{P}(g,f_{2})$ coincides with the $(\mathcal{E},\mathcal{M})$-factorization of $\mathcal{K}(g,L)$. In other words $\mathcal{P}(g,f_{2})\cong\mathcal{P}(g,L)$. Then, since $(a)$ is a pullback, $\mathcal{P}(g,f)$ is also the pullback of $\mathcal{P}(g,L)\rightarrowtail\mathcal{K}(A,L)$; thus $\mathcal{P}(g,f)\cong\mathcal{P}(g,f_{1})$ and $r\cong r_{1}$. Since $f_{1}$ is $(\lambda,\mathcal{E})$-pure this means that $r$ is in $\mathcal{E}$, as required. ∎
@@
Let $f=f_{2}f_{1}$ be $(\lambda,\mathcal{E})$-pure. Then there is a morphism $s\colon\mathcal{P}(g,f_{1})\to\mathcal{P}(g,f)$ induced by the universal property of the pullback defining $\mathcal{P}(g,f)$; this $s$ makes the diagram below commute.
@@
Since the maps into $\mathcal{K}(A,K)$ are in $\mathcal{M}$, so is $s$. Moreover, since the factorization system is proper and $r$ is in $\mathcal{E}$, also $s$ is in $\mathcal{E}$. Then $r$ is an isomorphism and $r_{1}$ is in $\mathcal{E}$; thus $f_{1}$ is $(\lambda,\mathcal{E})$-pure. ∎

### v3

@@ 1. Introduction
The concept of purity originated in the theory of abelian groups ([23]) and it is now well established both for modules ([22]) and in model theory in general ([29]). The monograph [2] made purity a key concept of the theory of accessible categories. For instance, given a locally $\lambda$-presentable category $\mathcal{K}$, it makes possible to characterize classes of objects injective with respect to a family of morphisms whose domains and codomains are $\lambda$-presentable, as classes closed under products, $\lambda$-directed colimits, and $\lambda$-pure subobjects (see [27]). Model-theoretically, this is a characterization of classes of models axiomatizable by regular theories in the logic $L_{\lambda\lambda}$ of $\lambda$-ary formulas.
@@
where $A$ and $B$ are $\lambda$-presentable, there is a morphism $t\colon B\to K$, such that $tg=u$ (see [2, 2.27]). Equivalently, given $u\colon A\to K$ then the existence of $v\colon B\to L$ such that $vg=fu$, implies the existence of $t\colon B\to K$ with $tg=u$. Since the existence of a morphism can be reformulated using surjectivity in the category $\operatorname{\bf Set}$ of sets, ordinary purity stems from the class $\mathcal{E}$ of surjections in $\operatorname{\bf Set}$.
Building on [15], our enriched purity will depend on a chosen class $\mathcal{E}$ of morphisms in $\mathcal{V}$ which, in addition, arises as the left class of a factorization system. In metric-enriched categories these are the dense morphisms.
@@
For $\mathcal{V}=\operatorname{\bf Met},\operatorname{\bf CMet},\operatorname{\bf Ban}$ or $\omega$-$\operatorname{\bf CPO}$ with $\mathcal{E}$ consisting of the dense maps, then $\mathcal{G}=\{I\}$ is the singleton determined by the monoidal unit (so that powers by $\mathcal{G}$ are trivial); instead, when $\mathcal{V}=\mathbf{DGAb}$ and $\mathcal{E}$ is formed by the regular epimorphisms, then $\mathcal{G}=\{P_{n}\}_{n\in\mathbb{Z}}$ is given by the suspensions and de-suspensions of the mapping cone of the unit (Example 8.4).
@@
While in this paper we already answer important questions about the relevance of enriched purity, there are still many open problems that wait to be addressed. For instance, when working with categories of structures of some relational language, ordinary pure morphisms correspond to the model theoretical notion of morphisms elementary with respect to positive-primitive formulas [2, 5.15]. A first glance at an enriched counterpart is given in Appendix A. However, a general answer cannot be provided at this stage for the simple reason that enriched counterparts of these logical concepts have not yet been defined. We believe that a thorough study of enriched purity will be essential for the development of enriched notions of languages and models, thus covering an outstanding gap in the literature.
@@ 2. Background
Consider a $\mathcal{V}$-category $\mathcal{K}$ and an ordinary functor $H\colon\mathcal{C}\to\mathcal{K}_{0}$ into the underlying ordinary category of $\mathcal{K}$; this corresponds to a $\mathcal{V}$-functor $\mathcal{C}_{\mathcal{V}}\to\mathcal{K}$ from the free $\mathcal{V}$-category on $\mathcal{C}$. When $\mathcal{C}$ is small, the (conical) limit of $H$ in $\mathcal{K}$ is the data of an object $\lim H\in\mathcal{K}$ together with isomorphisms in $\mathcal{V}$
@@
that are $\mathcal{V}$-natural in $K\in\mathcal{K}$, where on the right-hand-side we have the ordinary limit in $\mathcal{V}_{0}$. It follows that $\lim H$, when it exists, is also the limit of $H$ in $\mathcal{K}_{0}$; moreover $\lim H$ corresponds to the limit of $H$ weighted by $\Delta I\colon\mathcal{C}_{\mathcal{V}}\to\mathcal{V}$ (see [12, 3.51]).
@@
Local presentability will also be a key concept in the following sections. For ordinary locally presentable categories our standard reference is [2]. In the enriched setting the concept was introduced by Kelly [13]: given a $\mathcal{V}$-category $\mathcal{K}$ with (conical) $\lambda$-filtered colimits, we say that $K\in\mathcal{K}$ is $\lambda$-presentable (in the enriched sense) if $\mathcal{K}(K,-)\colon\mathcal{K}\to\mathcal{V}$ preserves $\lambda$-filtered colimits. We denote by $\mathcal{K}_{\lambda}$ the full subcategory of $\mathcal{K}$ spanned by the $\lambda$-presentable objects; this is closed in $\mathcal{K}$ under $\lambda$-small conical colimits and copowers by objects in $\mathcal{V}_{\lambda}$.
@@
For the remainder of this paper we shall assume our base of enrichment $\mathcal{V}$ to be locally $\lambda$-presentable as a closed category [13], meaning that $\mathcal{V}_{0}$ is locally $\lambda$-presentable, $I$ is $\lambda$-presentable and the $\lambda$-presentable are closed under tensor product (in particular $\mathcal{V}$ is then locally $\lambda$-presentable as a $\mathcal{V}$-category). In that context, many results from the ordinary theory of local presentability extend to the enriched context; see [13] for details.
Finally, (orthogonal) factorization systems will be central in the definition of our enriched notions of purity. Following [10] we will say that a factorization system $(\mathcal{E},\mathcal{M})$ on a category $\mathcal{K}$ is proper if every element of $\mathcal{E}$ is an epimorphism and every element of $\mathcal{M}$ a monomorphism. By [10, 2.1.4], when the factorization system is proper its is true that whenever a composite $fg\in\mathcal{E}$ then also $f\in\mathcal{E}$; we will make use of this property quite often.
@@ 3. Enriched purity
We fix $\mathcal{V}$ to be locally $\lambda$-presentable as a closed category and $(\mathcal{E},\mathcal{M})$ to be factorization system on $\mathcal{V}_{0}$. In this section we introduce the first of two concepts of enriched purity and prove several properties that this satisfies.
@@
of the map $\mathcal{K}(g,L)$ in $\mathcal{V}$. Let then $\mathcal{P}(g,f)$ be the pullback of $\mathcal{P}(g,L)$ along $\mathcal{K}(A,f)$, together with the induced map $r\colon\mathcal{K}(B,K)\to\mathcal{P}(g,f)$ as depicted below.
@@
We say that $f\colon K\to L$ is $\mathcal{E}$-pure with respect to $g$ if the map $r$ above lies in $\mathcal{E}$. We say that $f$ is $(\lambda,\mathcal{E})$-pure if it is $\mathcal{E}$-pure with respect to every $g\colon A\to B$ with $A$ and $B$ $\lambda$-presentable.
Remark 3.3 below gives a possibly more intuitive interpretation of the notion just introduced. In (1) of Example 3.5 below we show that, over $\operatorname{\bf Set}$, we get the standard notion of purity. The reader can also look at a model theoretic characterization of purity from Appendix A in terms of elementary morphisms. We will spell-out in Sections 6, 7, and 8 what purity looks like for the bases of enrichment we are most interested in.
@@
Since the elements of $\mathcal{M}$ are stable under pullback, the map $r$ above lies in $\mathcal{E}$ if and only if $\mathcal{P}(g,f)$ is the $(\mathcal{E},\mathcal{M})$-factorization of $\mathcal{K}(g,K)$. Thus $f\colon K\to L$ is $(\lambda,\mathcal{E})$-pure if and only if:
for every $g\colon A\to B$, with $A$ and $B$ $\lambda$-presentable,
the $(\mathcal{E},\mathcal{M})$-factorization of $\mathcal{K}(g,K)$ is obtained by pulling back the
$(\mathcal{E},\mathcal{M})$-factorization of $\mathcal{K}(g,L)$ along $\mathcal{K}(A,f)$.
Remark 3.4.
Following [25], a morphism $f\colon V\to W$ in $\mathcal{V}$ is called surjective if $\mathcal{V}_{0}(I,f)$ is a surjection of sets. A map is called injective if it satisfies the unique right lifting property with respect to surjections. Quite often, (surjective, injective) forms a proper factorization system on $\mathcal{V}$ (see [25]).
@@
    Consider $\mathcal{V}$ such that (surjective, injective) is a factorization system, and note that by [25, 3.4], injective maps are monomorphisms. Under Notation 3.1, it follows that $\mathcal{V}_{0}(I,\mathcal{P}(g,L))$ is the (epi, mono) factorization of $\mathcal{K}_{0}(g,L)$, so that
@@
    Given $\mathcal{V}$, consider the factorization system $(\mathcal{E},\mathcal{M})=(\mathcal{V}^{\to},\mathcal{V}^{\cong})$, where $\mathcal{V}^{\cong}$ is the class of all isomorphisms and $\mathcal{V}^{\to}$ the class of all morphisms. Then, every morphism $f$ is $\mathcal{V}^{\to}$-pure with respect to any $g\colon A\to B$; indeed, the map $r\colon\mathcal{K}(B,K)\to\mathcal{P}(g,f)$ of Notation 3.1 always lies in $\mathcal{V}^{\to}$.
@@
    Consider instead the factorization system $(\mathcal{E},\mathcal{M})=(\mathcal{V}^{\cong},\mathcal{V}^{\to})$ on $\mathcal{V}$, so that $\mathcal{E}$ is the class of all isomorphisms. Given $f\colon K\to L$ and $g\colon A\to B$, it follows that $\mathcal{P}(g,L)\cong\mathcal{K}(B,L)$ and therefore
    $\mathcal{P}(g,f)\cong\mathcal{K}^{\to}(g,f)$
    is the hom-object in the $\mathcal{V}$-category of arrows $\mathcal{K}^{\to}$. Then $f$ is $\mathcal{V}^{\cong}$-pure with respect to $g$ if and only if
    $\mathcal{K}(B,K)\cong\mathcal{K}^{\to}(g,f);$
    looking at the underlying categories, this implies in particular that for every pair $u\colon A\to K$ and $v\colon B\to L$, such that $fu=vg$, there is a unique $t\colon B\to K$ such that $ft=v$ and $tg=u$.
    For $\mathcal{K}$ locally $\lambda$-presentable, $(\lambda,\mathcal{V}^{\cong})$-pure morphisms then coincide with the isomorphisms. On the one hand, isomorphisms are always $\mathcal{V}^{\cong}$-pure with respect to any morphism. Conversely, assume that $f$ is $(\lambda,\mathcal{V}^{\cong})$-pure, and write it as a $\lambda$-directed colimit of morphisms $f_{i}\colon K_{i}\to L_{i}$ between $\lambda$-presentable objects (see [2, 1.55(1)]); then for any $i$ we can consider the square
[…]
    where $(u_{i},v_{i})$ is part of the colimiting cocone of $f$. By $\mathcal{V}^{\cong}$-purity of $f$, for any $i$ there exists a unique $t_{i}\colon L_{i}\to K$ such that $ft_{i}=v_{i}$ and $t_{i}f_{i}=u_{i}$. It is easy to see that the $(t_{i},v_{i})$ also define a colimiting cocone for $f$ in $\mathcal{V}^{\to}$. Therefore $f$ is a colimit of the isomorphisms $1_{L_{i}}\colon L_{i}\to L_{i}$, and thus $f$ itself is an isomorphism.
@@
    Let $\operatorname{\bf CMet}$ be the category of complete metric spaces and nonexpanding maps, where we allow distances $\infty$. Here, we have the proper factorization system (dense, isometry) — see [3, 3.16(2)].
    Under this factorization system, given $g$ and $f$ as above, $\mathcal{P}(f,g)$ is the sub-metric space of $\mathcal{K}(A,K)$ consisting of those $u\colon A\to K$ such that
    for every $\varepsilon>0$ there exists $v\colon B\to L$ for which
    $fu\sim_{\varepsilon}vg$ (that is, $d(fu,vg)\leq\varepsilon$)
    Then $f$ is $\mathcal{E}$-pure with respect to $g$ if and only if for any $u\in\mathcal{P}(f,g)$ as above and for any $\varepsilon>0$, there exists $t\colon B\to K$ such that $tg\sim_{\varepsilon}u$.
    A priori this doesn’t correspond to a notion of purity studied for metric-enriched categories. However, we will see in the more general setting of Section 6, that this is an equivalent reformulation of the notions considered in [28].
@@
    In the category $\operatorname{\bf Met}$ of metric spaces and nonexpanding maps, we have two factorization systems: (dense, closed isometry) and (surjective, isometry) — see [3, 3.16]. The first one yields the same concept of purity described in (4) above, while the second one gives the usual (unenriched) concept of purity, as in (1).
    Observe that surjections in $\operatorname{\bf CMet}$ are the usual surjective maps but (surjective, injective), as defined in Remark 3.4, does not form a factorization system. For instance the morphism $\mathbb{N}\to\{\frac{1}{n}|n\in\mathbb{N}\}\cup\{0\}$ (sending $n$ to $\frac{1}{n}$), where the domain is discrete and the codomain has the usual metric, does not factorize.
@@
    Let $\mathcal{V}=\mathbf{Gra}$ be the cartesian closed category of directed graphs (possibly) without loops. Consider the (regular epi, mono) factorization system. Given $f\colon K\to L$, this is pure in the enriched sense if for any morphism $g\colon A\to B$ between finitely presentable objects, and any vertices $u\in\mathcal{K}(A,K)$ and $v\in\mathcal{K}(B,L)$ for which $vg=fu$, there exists a vertex $t\in\mathcal{K}(B,K)$ such that $tg=u$. Since not every vertex is a morphism in the underlying category (only the vertices with a loop are such), the notion of enriched purity is not the ordinary one on the underlying category; this is essentially because the unit is not regular projective.
@@
    Let $\mathcal{V}=\mathbf{Ban}$ be the category of Banach spaces over $\mathbb{C}$ with linear maps of norm at most $1$. This is symmetric monoidal closed with $\otimes$ given by the projective tensor product, and internal hom $[K,L]$ is the space consisting of all bounded linear operators from $K$ to $L$ (see [5, 6.1.9h]). Note that the monoidal unit $\mathbb{C}$ is a strong generator: $\mathbf{Ban}(\mathbb{C},-)$ is the unit-ball functor, and an operator is an isomorphism in $\mathbf{Ban}$ if and only if it is a bijection on the unit balls.
    Here, we can consider the (epi, strong mono) factorization system, which by [21, 1.15 & 2.3] coincides with the (dense, isometry) factorization system; this induces a notion of purity similar to that of $\mathbf{Met}$ — see in particular Examples 3.22 and 4.19.
    We can also consider (strong epi, mono) factorization system. It follows from [21, 2.7 & 2.10] that regular and strong epimorphisms in $\mathbf{Ban}$ coincide. Moreover, they are just the surjective operators, or equivalently, the operators $T$ for which $\mathbf{Ban}_{0}(\mathbb{C},T)$ is surjective ([21] shows that strong/regular epimorphisms are surjective, but since homming-out of a strong generator reflects strong epimorphisms, then also every surjective operator is a strong epimorphism). Since $I=\mathbb{C}$, these coincide with the surjections from Remark 3.4. Thus the notion of purity corresponding to this factorization system coincides with the ordinary one on the underlying category.
@@
the squares $(a)$, $(b)$, and $(a+b)$ are pullbacks. By $\mathcal{E}$-purity of $f_{2}$, the map $r_{2}$ lies in $\mathcal{E}$ and thus $\mathcal{P}(g,f_{2})$ coincides with the $(\mathcal{E},\mathcal{M})$-factorization of $\mathcal{K}(g,L)$. In other words $\mathcal{P}(g,f_{2})\cong\mathcal{P}(g,L)$. Then, since $(a)$ is a pullback, $\mathcal{P}(g,f)$ is also the pullback of $\mathcal{P}(g,L)\rightarrowtail\mathcal{K}(A,L)$ along $\mathcal{K}(A,f_{1})$; thus $\mathcal{P}(g,f)\cong\mathcal{P}(g,f_{1})$ and $r\cong r_{1}$. Since $f_{1}$ is $(\lambda,\mathcal{E})$-pure this means that $r$ is in $\mathcal{E}$, as required. ∎
@@
Let $f=f_{2}f_{1}$ be $(\lambda,\mathcal{E})$-pure. Then there is a morphism $s\colon\mathcal{P}(g,f_{1})\to\mathcal{P}(g,f)$ induced by the universal property of the pullback defining $\mathcal{P}(g,f)$; this $s$ makes the diagram below commute,
@@
where $r_{1}$ denotes the dashed map from Notation 3.1 corresponding to $f_{1}$. Since the morphisms into $\mathcal{K}(A,K)$ are in $\mathcal{M}$, so is $s$. Moreover, since the factorization system is proper and $r$ is in $\mathcal{E}$, also $s$ is in $\mathcal{E}$. Therefore $r$ is an isomorphism and $r_{1}$ is in $\mathcal{E}$; thus $f_{1}$ is $(\lambda,\mathcal{E})$-pure. ∎